Probability
Get insights from 70 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability
- 70
Questions
- 0
Discussions
- 7
Active Users
- 0
Followers
Have a question about Probability? Ask our experts
Keep it short & simple. Type complete word. Avoid abusive language.
49. Out of 100 students, two sections of 40 and 60 are formed. If you and your friendare among the 100 students, what is the probability that
(a) You both enter the same section?
(b) You both enter the different sections?
49. Out of 100 students, two sections of 40 and 60 are formed. If you and your friendare among the 100 students, what is the probability that
(a) You both enter the same section?
(b) You both enter the different sections?
49. Here, out of 100 students, first section has 40 students and the rest I e, 60 students enters in second section.
As me and my friend are among the 100 students.
The no. of ways of selecting 2 students from the 100 students
= 100C2
(a) When both enters first section if 2 of us are among the 40 stud
48. In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy (a) One ticket (b) Two tickets (c) 10 tickets.
48. In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy (a) One ticket (b) Two tickets (c) 10 tickets.
48. Total no. of ticket for lottery sold = 10, 000
No. of ticket that are awarded prize = 10
So, no. of ticket that are not awarded prize = 10.000 - 10
= 9990
(a) Now, probability of met getting a prize if we buy one ticket

47. A die has two faces each with number ‘1’, three faces each with number ‘2’ and one face with number ‘3’. If die is rolled once, determine
(i) P(2) (ii) P(1 or 3) (iii) P(not 3)
47. A die has two faces each with number ‘1’, three faces each with number ‘2’ and one face with number ‘3’. If die is rolled once, determine
(i) P(2) (ii) P(1 or 3) (iii) P(not 3)
47. Since the die has two faces each mark 1, three faces each marks 2 and one face mark 3.
The possible sample space of outcome is.
S = {1, 2, 3} so, n (s) = 6
(i) P (2)
(ii) P (1 or 3) = P (1) + P (3) =
(iii) P (not 3) = 1 P (3) = 1 - =
46. 4 cards are drawn from a well – shuffled deck of 52 cards. What is the probability of obtaining 3 diamonds and one spade?
46. 4 cards are drawn from a well – shuffled deck of 52 cards. What is the probability of obtaining 3 diamonds and one spade?
46. Total number of ways of drawing 4 cards from a duck of 52 cards, n (s) = 52C4
Total no. of diamond cards = 13
Similarly, total. no. of spades cards = 13

45. A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that
(i) All will be blue? (ii) Atleast one will be green?
45. A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that
(i) All will be blue? (ii) Atleast one will be green?
45. Given,
No. of red marbles= 10
No. of blue marbles = 20
No. of green marbles = 30.
So, total no. of marbles = 10 + 20 + 30 = 60
Now, we are to select 5 marbles from the given 60 marbles.
So the sample space is:
n (s) = 60c5.


44. In a class of 60 students, 30 opted for NCC, 32 opted for NSS and 24 opted fo rboth NCC and NSS. If one of these students is selected at random, find the probability that
(i) The student opted for NCC or NSS.
(ii) The student has opted neither NCC nor NSS.
(iii) The student has opted NSS but not NCC.
44. In a class of 60 students, 30 opted for NCC, 32 opted for NSS and 24 opted fo rboth NCC and NSS. If one of these students is selected at random, find the probability that
(i) The student opted for NCC or NSS.
(ii) The student has opted neither NCC nor NSS.
(iii) The student has opted NSS but not NCC.
44. Given that, total number of student, n (S) = 60
Let A: student opted for NCC
n (A) = 30
B: student opted for NSS
n (B) = 32
And student who opted both NCC and NSS, n (A∩B) = 24
(i) Probability that student opted for NCC or NSS,
P (A∪B) = P (A) + P (B) – P (A∩B)
(ii) Probability that student opted nei
43. The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?
43. The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?
43. Let A: student passing in Hindi
B: student passing in English
Given, P (B) = 0.75
P (A∩B) = 0.5, passing both subject
And P (A'∩B') = 0.1, i.e., passing neither subject
P (A∪B)' = 0.1
1 – P (A∪B) = 0.1
P (A∪B) = 1 – 0.1
P (A∪B) = 0.9
Hence, P (A∪B) = P (A) + P (B) – P (A∩B)
P (A) = P (A∪B) + P (A∩B) – P (
42. In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing atleast one of them is 0.95. What is the probability of passing both?
42. In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing atleast one of them is 0.95. What is the probability of passing both?
42. Let A: Student passes 1st examination
So, P (A) = 0.8
And B: Student passes 2nd examination
So, P (B) = 0.7
Also probability of passing at least one examination is P (A∪B) = 0.95
Therefore, P (A∪B) = P (A) + P (B) – P (A∩B)
0.95 = 0.8 + 0.7 – P (A∩B)
P (A∩B) = 0.8 + 0.7 – 0.95
P (A∩B) = 0.55
Hence,
41. In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.
41. In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.
41. Given that, 40% study Mathematics, 30% study Biology and 10% study both Mathematics and Biology.
Let A: Students study Mathematics.
P (A) = 40% = .
Let B: Students study Biology.
P (B) = 30% = .
So, P (A∩B) i.e. probability of student studying both Mathematics and Biology is
P
40. A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16.Determine
(i) P(not A), (ii) P(not B) and (iii) P(A or B)
40. A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16.Determine
(i) P(not A), (ii) P(not B) and (iii) P(A or B)
40. Given P (A) = 0.42
P (B) = 0.48
P (A∩B) = 0.16
(i) P (not A) = P (A') = 1 – P (A) = 1 – 0.42 = 0.58
(ii) P (not B) = P (B') = 1 – P (B) = 1 – 0.48 = 0.52
(iii) P (A or B) = P (A∪B) = P (A) + P (B) – P (A∩B)
= 0.42 + 0.48 – 0.16
= 0.74